OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. 70 72 111 114

OpenStudy (raden):

first, a27 = 66 means a1 + (27-1)d = 66 -12 + 26d = 66 solve for d

OpenStudy (anonymous):

3

OpenStudy (raden):

yep.. correct

OpenStudy (anonymous):

k

OpenStudy (raden):

now, to get a42 use the formula : an = a1 + (n-1)d a42 = -12 + (42-1)3 a42 = -12 + 41(3) = ... ?

OpenStudy (raden):

got it ?

OpenStudy (raden):

what's 41 times 3 ..... then subtracted by 12

OpenStudy (anonymous):

111, !

OpenStudy (raden):

cool :D